OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y - x2 = 11x + 30?

OpenStudy (anonymous):

start with \(y=x^2+11x+30\) set \(y=0\) and solve \[x^2+11x+30=0\]

OpenStudy (anonymous):

this one factors, so it should not be hard to find the zeros

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw